# Vi-LAD: Vision-Language Attention Distillation for Socially-Aware Robot Navigation in Dynamic Environments

### v1

@@ VI Conclusion, Limitations & Future work
We introduced Vi-LAD, a novel Vision-Language Attention Distillation method that enables socially compliant and real-time robotic navigation by distilling knowledge from large Vision-Language Models (VLMs) into a lightweight transformer-based model. By leveraging attention map-level distillation, Vi-LAD effectively integrates social navigation reasoning from both a pre-trained vision-action model and a VLM, ensuring efficient and adaptive motion planning. Our approach demonstrates significant improvements over SOTA methods in real-world experiments on a Husky robot, with higher success rates and trajectories that align closely with human teleoperated actions. However, while our attention-guided MPC planner enhances social compliance, it may require further tuning for highly crowded settings. Future work will explore pre-trained models with additional modalities, such as depth and LiDAR, to enhance spatial awareness and robustness. We also plan to investigate online adaptation strategies for real-time refinement, and extensions to long-horizon navigation in complex environments.

### v2

@@ VI Conclusion, Limitations & Future work
We introduced ViLAM, a novel Vision-Language Attention Distillation method that enables socially compliant and real-time robotic navigation by distilling knowledge from large Vision-Language Models (VLMs) into a lightweight transformer-based model. By leveraging attention map-level distillation, ViLAM effectively integrates social navigation reasoning from both a pre-trained vision-action model and a VLM, which ensures efficient and adaptive motion planning. Our approach demonstrates significant improvements over SOTA methods in real-world experiments on a Husky robot, with higher success rates and trajectories that align closely with human teleoperated actions. However, as the distilled social priors are learned offline, our attention-guided planner may require further tuning in highly crowded or previously unseen settings. Future work will explore pre-trained models with additional modalities, such as depth and LiDAR, to enhance spatial awareness and robustness. We also plan to investigate online adaptation strategies for real-time refinement, and extensions to long-horizon navigation in complex environments.